I am having a hard time understanding the following equation;
(-2)(4)+5-6-[-2]
________________= ?
(-2)[9+(-5)]
What are the [] brackets used for in the equation? This is one equation.
thanks........

Answer by checkley75(3666)

() & [] are the same.
(-2)(4)=-8
+5-6=-1
1[-2]=2
---------------------------
(-2)[9-5]=(-2)(4)=-8
now combine these terms thus:
(-8-1+2)/-8
-7/-8=7/8 is the answer.